OpenStudy (anonymous):

I'm confused as to how to graph this. I'm suppose to find at least three ordered pairs. It say's: When the coefficient of x is a fraction, choose values of x that will simplify the calculations.

OpenStudy (anonymous):

Is this referring to your post of f(x)=-1/2+3? If so, then to plot points, you should choose values of x that are multiples of 2 so that the fraction simplifies. For example: f(2)=-1/2(2)+3=-1+3=2 f(6)=-1/2(6)+3=-3+3=0 f(10)=-1/2(10)=-5+3=-2 You can pick any numbers you want, but if you choose x-values that are odd (ex: 3) the output will be a fraction (ex: f(3)=-1/2(3)+3=-3/2+3=-3/2+6/2=3/2).